# SemCSE-Multi: Multifaceted and Decodable Embeddings for Aspect-Specific and Interpretable Scientific Domain Mapping

### v1

@@ 1 Introduction
Beyond citation networks for visual exploration van Eck and Waltman (2010, 2014), embedding models have been proposed as a promising remedy Cohan et al. (2020); Ostendorff et al. (2022b); Singh et al. (2023); Brinner and Zarriess (2025). By mapping scientific abstracts into vector representations, these models enable efficient similarity assessment and support large-scale search as well as semantic clustering and visualization of a scientific domain.
@@
In this work, we mainly consider the case of invasion biology, a scientific field investigating the introduction of non-native species into new ecosystems. Hypotheses about general causal relationships (e.g., the influence of ecosystem properties on the likelihood of species invasions) play a central role in this domain, so that a researcher might regard studies as related if they address the same hypothesis, regardless of other factors like the particular species or ecosystem under consideration.
Another researcher, however, may instead be interested in specific species, without any requirements concerning the underlying hypothesis, thus regarding vastly different studies as highly related. For this reason, a broadly useful embedding space for literature retrieval or semantic visualization (see Figure 1) must support aspect-specific and user-adaptable similarity assessments, rather than relying on a single, pre-defined notion of similarity.
To address both limitations, we propose a comprehensive framework for training a multifaceted embedding model that encodes distinct, user-specified aspects of scientific papers into separate embeddings, thus enabling precise and aspect-specific similarity assessment. In addition, the framework supports the reconstruction of natural language descriptions from embeddings, thus providing direct interpretability of the embedding space. Combined with established techniques for visualizing high-dimensional embedding spaces, our approach therefore enables the creation of controllable, interpretable, and interactive visualizations of scientific domains.
At a high level, our framework consists of three core components. First, building on recent advances in training semantically focused embedding models via LLM-generated summaries Brinner and Zarriess (2025), we train multiple aspect-specific embedding models. Each model focuses solely on one distinct aspect, and is trained by generating multiple aspect-specific summaries for each scientific abstract and optimizing the model to place them in close proximity within the embedding space (Section 3).
@@
Third, we design a decoding mechanism that leverages an LLM to reconstruct natural language representations of embeddings, thus substantially enhancing interpretability of the embedding space (Section 5). We further extend this interpretability to low-dimensional visualizations, allowing for the decoding of previously unoccupied regions of the embedding space into meaningful textual descriptions (Section 6).
We evaluate our framework primarily in the domain of invasion biology, for which we had guidance from domain experts. To highlight its broader applicability, we further present a smaller-scale evaluation in the medical domain in Appendix A.8.
@@ 2 Related Work
Visualizing Scientific Domains has been approached through a variety of methods, with the most prominent being the visualization of citation networks. In these approaches, papers that cite one another are treated as related and clustered accordingly van Eck and Waltman (2010, 2014). Building on this idea, various embedding models Cohan et al. (2020); Ostendorff et al. (2022b); Singh et al. (2023) have been trained on citation information to place papers that cite each other in close proximity within an embedding space, thereby enabling similarity comparisons without requiring direct access to the citation network. However, relying on citation information as a proxy for similarity provides no insight or control into which aspects of the papers are being considered for assessing similarity.
@@
A common limitation across both general and multifaceted embedding models is the lack of user control in specifying which aspects of similarity should be taken into account, which is only possible within methods that rely on large supervised datasets, which are costly to acquire. Our work addresses this limitation by introducing a fully unsupervised method for constructing multifaceted embeddings that explicitly encode the aspects most relevant to potential users.
@@ 3 Training Aspect Embedding Models
The first step in our embedding framework is the creation of several distinct embedding models, each capable of mapping summaries of specific aspects into a dedicated semantic embedding space.
3.1 Motivation
Our approach is motivated by the observation that multiple aspects can be relevant when assessing the similarity of studies within a scientific domain, and that the relative importance of these aspects may vary depending on the user’s goals. If, for example, a database of scientific papers shall be visualized such that distances in the embedding space reflect similarity (see Figure 1), users may therefore wish to configure the visualization so that only similarity with respect to certain aspects is taken into account, while disregarding other, less relevant ones.
As a case study, we focus on the domain of invasion biology (with experiments in the medical domain being reported in Appendix A.8), which investigates the introduction of non-native species into new ecosystems and their ecological, economic, and societal impacts. In this domain, the aspects that determine similarity between studies vary substantially depending on the research perspective. For example, one user may be primarily interested in studies addressing the same hypothesis, while another may focus on the same species or ecosystem type (see Section 1). Beyond these primary factors, additional aspects such as the precise research question, specific recommendations, or the methodology employed (e.g., field surveys versus laboratory trials) can serve as important auxiliary indicators of similarity, helping to shape a more fine-grained and meaningful structure of the embedding space. A detailed discussion of these relevant aspects is provided in Appendix A.1.
@@ 3.2 Dataset Creation
The foundation of our framework is a dataset of scientific abstracts from the target domain. For this purpose, we leverage a corpus of 37,786 papers in the field of invasion biology, compiled by Brinner et al. (2025). To train the aspect-specific embedding models, we then require multiple summarizing sentences for each abstract and aspect, that summarize the aspect-specific information of that paper for just the aspect of interest in isolation, while excluding information about other factors.
To generate these summaries, we use Mistral 3.1 Small Mistral AI (2025), prompting it four times per abstract to summarize the designated aspect. Due to stochasticity in autoregressive generation, the resulting summaries differ slightly in phrasing but remain highly semantically aligned because of the shared generation context. If the aspect does not apply to a given study, the model is instructed to return “Not applicable”. Further details on important insights, exemplary prompts and dataset sizes for each factor are reported in Appendix A.2, with exemplary summaries being displayed in Figure 3.
@@ 3.3 Model Training
Hypotheses Ecosystems Res. Questions Species Methodology Recommendation
@@
For each aspect separately, we train a dedicated embedding model following the approach of Brinner and Zarriess (2025). Specifically, we apply a contrastive learning objective that encourages the model to place two summarizing sentences from the same abstract close together in the embedding space, while pushing apart embeddings of unrelated sentences (i.e., summarizing the same aspect for other abstracts):
$\mathcal{L}=-\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\log\frac{{\rm e}^{\text{sim}(e_{i},e_{i}^{+})/\tau}}{\sum_{j}{\rm e}^{\text{sim}(e_{i},e_{j}^{+})/\tau}}$
Here, the loss is computed over a batch $\mathcal{B}$ consisting of $|\mathcal{B}|$ pairs of related samples, represented by their embeddings $(e_{i},e_{i}^{+})$. As the similarity function, we use cosine similarity, with $\tau$ being a temperature parameter.
As the base encoder, we employ SciDeBERTa v2 Kim et al. (2023), mapping the CLS token embedding down to 150150 dimensions with an additional linear projection head. Additional implementation and training details are provided in Appendix A.3.
@@ 3.4 Evaluation Setup
Since the aspect-specific embedding models are merely a intermediate result, we perform a small-scale evaluation in a retrieval setting. To this end, we are using a separate test set of 954 scientific abstracts from the field of invasion biology Brinner et al. (2022), for which aspect-specific summaries are generated in the same manner as for training.
@@ 3.5 Results
To underscore the aspect-specific nature of the models, we conduct an ablation study in which, for each aspect, retrieval is performed using all models trained on the other aspects. In this setting, performance drops substantially, thus providing clear evidence that each model learns to capture information specific to its designated aspect. We also evaluate SemCSE as baseline, which cannot match the aspect-specific models as well.
@@ 4.1 Model Training
Concretely, we employ a pretrained SciDeBERTa v2 encoder Kim et al. (2023) with several linear projection layers - one for each aspect - that map the 768-dimensional CLS token embedding into the 150-dimensional aspect-specific embeddings. The model is trained using an L2 loss between the predicted and target embeddings. Further experimental details are provided in Appendix A.4.
@@ 4.2 Evaluation Setup
    The test set of 954 invasion biology abstracts is annotated with labels indicating which of the ten most prominent hypotheses each abstract addresses. A well-performing embedding model should, for the hypothesis aspect, place abstracts that address the same hypothesis close together in the embedding space.
@@
In the latter case, Qwen3-30B-A3B Yang et al. (2025) is prompted to assign integer similarity scores to pairs of abstracts, strictly with respect to a given aspect and based on precise category definitions (Appendix A.5). While these assessments are likely of lower quality than expert annotations, strong correlations with embedding-based similarities nevertheless indicate that the embeddings capture aspect-specific information, whereas low alignment with embeddings for a different aspect suggests that the unified model can effectively disentangle and isolate aspect-specific information.
@@ 4.3 Results
Results for this evaluation are displayed in Table 2. The upper part reports the correlations between the aspect-specific embeddings created by SemCSE-Multi and the different ground-truth similarity assessments. For completeness, we include the human-annotated hypothesis labels as predictors, allowing us to assess their correlation with the LLM-produced similarity scores.
@@
    The human-annotated hypothesis labels strongly correlate with the LLM-based similarities for the hypothesis aspect, thus confirming the validity of the LLM assessments.
@@
    Aspect-embeddings for the hypothesis aspect correlate well with the human-annotated hypothesis labels, thus demonstrating that this factor is effectively represented in the embedding space.
@@
The substantially lower correlations observed in the remaining off-diagonal cells provide strong evidence that the aspect-specific embeddings are indeed isolating the intended factors. In consequence, this disentanglement enables controlled similarity assessments adhering to specific user needs.
One notable exception is the recommendation aspect, where embeddings only show weak alignment with the LLM-based similarities. This is likely due to the scarcity of explicit recommendations in scientific abstracts, which reduces the number of training samples. Moreover, in cases where recommendations are absent or solely hinted at, the LLM occasionally produces summaries that are not sufficiently supported by the abstract, thereby degrading the training data quality. Naturally, this problem is magnified for aspects with far more samples that not sufficiently address a specific aspect.
To contextualize the strength of the observed correlations, we further compare our results against several widely used scientific embedding models: SPECTER Cohan et al. (2020), SPECTER2 Proximity Singh et al. (2023), SciNCL Ostendorff et al. (2022b), SemCSE Brinner and Zarriess (2025), and the invasion biology-specific InvDef-DeBERTa Brinner et al. (2025). Across all aspects except recommendation, none of these models achieve correlations comparable to our aspect-specific embeddings. Furthermore, these baseline models cannot disentangle specific aspects, making them unsuitable for tasks requiring fine-grained similarity assessments. This is underscored by Figure 1, visualizing the clearly differing and semantically structured aspect-specific embedding spaces by SemCSE-Multi for the Hypothesis and Species aspects, as well as the substantially less distinctly structured space induced by the SciNCL model, which is not adaptable to any specific aspect. Together, these results provide clear evidence that a multifaceted embedding model offers distinct advantages both in terms of accuracy and in the ability to isolate aspect-specific information.
@@ 5 Embedding Decoding
In this section, we demonstrate that it is possible to use a decoder LLM to map aspect-specific embeddings back into the space of natural language descriptions. This capability enables direct interpretability of the embedding space and supports interactive exploration grounded in semantically meaningful explanations.
@@ 5.1 Model Training
To guide the decoding process, we additionally prepend five trainable, input-independent embeddings that function as general task descriptors, and - for the Llama model - append one more token to act as delimiter to the decoded text. During training, only these embeddings and the mapping network are optimized, while the parameters of the underlying LLM remain frozen. Using the dataset of summarizing sentences paired with their corresponding aspect-specific embeddings, the model is trained with a standard language modeling cross-entropy loss to reconstruct the original summary from its embedding representation.
We evaluate two LLM backbones of different sizes and training paradigms, with one being a purely autoregressive language model in the form of Llama 3 8B Grattafiori et al. (2024), and the other being a larger (24B parameters) chat model in the form of Mistral Small 3.1 Mistral AI (2025). Note that, for the chat model, the learned and predicted tokens are injected into the user prompt while retaining the standard chat formatting.
@@ 5.2 Evaluation Setup
Table 3: Perplexity results for Llama and Mistral decoders when decoding embeddings under normal, reconstructed and randomized settings. Lower perplexity indicates better performance.
We evaluate the decoding pipeline on the test set of 954 abstracts for which we predict aspect-specific embeddings with SemCSE-Multi. For a given aspect embedding, we then measure the perplexity of the target reconstruction (i.e., the up to four aspect-specific summarizing sentences) under the predicted distribution produced by the LLM. Low perplexity in this setting indicates that the decoder is able to recover the semantic content of the original summary with high probability.
To verify that the model does not merely generate generic aspect-related sentences while ignoring the actual embedding input, we additionally evaluate the perplexity when matching the embedding of one sample with the summaries of another (Shuffled). Further, for the pure language model (Llama), we evaluate the perplexity of the summaries without any embedding information (Unconditioned). If the trained decoding process indeed reconstructs the semantic meaning of the embeddings, perplexity of the reconstruction should be substantially lower than that in these control experiments.

### v2

@@ 1 Introduction
Embedding models have been identified as a promising remedy: by mapping scientific abstracts into dense vector representations, these models enable efficient similarity assessment and support large-scale search as well as semantic clustering and visualization of a scientific domain Cohan et al. (2020); Ostendorff et al. (2022b); Singh et al. (2023); Brinner and Zarriess (2025).
@@
In this work, we mainly consider the scientific field of invasion biology, which investigates the introduction of non-native species into new ecosystems and their ecological, economic, and societal impacts. Hypotheses about general causal relationships (e.g., the influence of ecosystem properties on the likelihood of species invasions) play a central role in this domain, so that a researcher might regard studies as related if they address the same hypothesis, regardless of other factors like the particular species or ecosystem under consideration.
Another researcher, however, may instead be interested in specific classes of species or ecosystems, without any requirements on the underlying hypothesis, thus regarding vastly different studies as highly related. Beyond these primary factors, additional aspects such as the precise research question, specific recommendations, or the methodology employed can serve as important auxiliary indicators of similarity, helping to shape a more fine-grained and meaningful structure of the embedding space.
For this reason, a broadly useful embedding space that meets a user’s requirements for literature retrieval or semantic visualization (see Figure 1) must support aspect-specific and user-adaptable similarity assessments, rather than relying on a single, pre-defined notion of similarity.
To address both adaptability and interpretability, we therefore propose a comprehensive framework for training a multifaceted embedding model that encodes distinct, user-specified aspects of scientific papers into separate embeddings, thus enabling precise, aspect-specific similarity assessment. Further, the framework supports the reconstruction of natural language descriptions from embeddings, thus providing direct interpretability of the embedding space and leading to controllable, interpretable, and interactive visualizations of scientific domains.
At a high level, our framework consists of three core components. First, building on recent advances in training semantically focused embedding models via LLM-generated summaries (SemCSE, Brinner and Zarriess, 2025), we train multiple aspect-specific embedding models. Each model focuses solely on one distinct aspect, and is trained to place multiple LLM-generated, aspect-specific summaries for a scientific abstract in close proximity within the embedding space (Section 3).
@@
Third, we design a decoding mechanism that reconstructs natural language representations of given embeddings, thus substantially enhancing interpretability of the embedding space (Section 5). Section 6 further extends this interpretability to low-dimensional visualizations, allowing for the decoding of previously unoccupied regions of the visualization into meaningful textual descriptions.
We evaluate our framework primarily in the domain of invasion biology, for which we had guidance from domain experts. To highlight its broader applicability, we further present a smaller-scale evaluation in the medical domain in Appendix B.
@@ 2 Related Work
Visualizing Scientific Domains has been approached through a variety of methods. Citation networks treat papers that cite one another as related and cluster them accordingly van Eck and Waltman (2010, 2014), and various embedding models build on this idea by being trained on citation information to place papers that cite each other in close proximity within an embedding space, thereby enabling similarity comparisons without requiring direct access to the citation network Cohan et al. (2020); Ostendorff et al. (2022b); Singh et al. (2023). However, relying on citation information as a proxy for similarity provides no insight or control into which aspects of the papers are being considered for assessing similarity.
@@
A common limitation across both general and multifaceted embedding models is the lack of user control in specifying which aspects of similarity should be taken into account, which is only possible within methods that rely on large supervised datasets. Our work addresses this limitation by introducing a fully unsupervised method for constructing multifaceted embeddings that explicitly encode the aspects most relevant to potential users.
@@ 3 Training Aspect Embedding Models
The first step in our embedding framework is the creation of several distinct embedding models, each capable of mapping summaries of specific aspects into a dedicated semantic embedding space. The six aspects under consideration were motivated in Section 1 and are detailed in Appendix A.1.
@@ 3.1 Dataset Creation
The foundation of our framework is a dataset of scientific abstracts from the target domain. For this purpose, we leverage a corpus of 37,786 papers from the field of invasion biology, compiled by Brinner et al. (2025). To train the aspect-specific embedding models, we then require multiple summarizing sentences for each abstract and aspect, that summarize the aspect-specific information of that paper for just the aspect of interest in isolation, while excluding information about other factors.
To generate these summaries, we use Mistral 3.1 Small Mistral AI (2025), prompting it four times per abstract to summarize the designated aspect or reply “Not applicable” if the aspect does not apply to the study. Due to stochasticity in autoregressive generation, the resulting summaries differ slightly in phrasing but remain highly semantically aligned because of the shared generation context. Important insights, exemplary prompts and dataset sizes for each factor are reported in Appendix A.2, with exemplary summaries being displayed in Figure 3.
@@ 3.2 Model Training
Hypotheses Ecosystem Res. Question Species Methodology Recommendation
@@
For each aspect separately, we train a dedicated embedding model following the SemCSE approach of Brinner and Zarriess (2025). Specifically, we apply a contrastive loss that encourages the model to place two aspect summaries from the same abstract close together in the embedding space, while pushing apart embeddings of unrelated summaries (i.e., same aspect but different abstract):
$\mathcal{L}=-\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\log\frac{{\textrm{e}}^{\text{sim}(e_{i},e_{i}^{+})/\tau}}{\sum_{j}{\textrm{e}}^{\text{sim}(e_{i},e_{j}^{+})/\tau}}$
Here, the loss is computed over a batch $\mathcal{B}$ consisting of $|\mathcal{B}|$ pairs of related samples with embeddings $(e_{i},e_{i}^{+})$. As the similarity function, we use cosine similarity, with $\tau$ being a temperature parameter. As the base encoder, we employ SciDeBERTa Kim et al. (2023), mapping the CLS token embedding down to $150$ dimensions with an additional linear projection head. Additional implementation and training details are provided in Appendix A.3.
@@ 3.3 Evaluation Setup
Since the aspect-specific embedding models are merely an intermediate result, we perform a small-scale evaluation in a retrieval setting. To this end, we are using a separate test set of 954 invasion biology abstracts Brinner et al. (2022) with aspect-specific summaries (generated as for training).
@@ 3.4 Results
To underscore the aspect-specific nature of the models, we conduct an ablation study in which, for each aspect, retrieval is performed using all models trained on the other aspects. In this setting, performance drops substantially, thus providing clear evidence that each model learns to capture information specific to its designated aspect. We also evaluate SemCSE as baseline, which was trained on general scientific texts and cannot match the aspect-specific models as well, and the base SciDeBERTa decoder, which achieves the lowest scores.
@@ 4.1 Model Training
Concretely, we employ a pretrained SciDeBERTa encoder Kim et al. (2023) with several linear projection layers - one for each aspect - that map the 768-dimensional CLS token embedding into the 150-dimensional aspect-specific embeddings. The model is trained using an L2 loss between the predicted and target embeddings. Further experimental details are provided in Appendix A.4.
@@ 4.2 Evaluation Setup
    The test set of 954 invasion biology abstracts is annotated by domain experts with labels indicating which of the ten most prominent hypotheses each abstract addresses. Thus, for the hypothesis aspect, the model should place abstracts that address the same hypothesis in close proximity within the embedding space.
@@
In the latter case, Qwen3-30B-A3B Yang et al. (2025) is prompted to assign integer similarity scores to pairs of abstracts, strictly with respect to a given aspect and based on precise category definitions (Appendix A.5). While these assessments are less reliable than the expert annotations, strong correlations with embedding-based similarities nevertheless indicate that the embeddings capture aspect-specific information. Further, low alignment of embedding similarity for one aspect with LLM-based scores for a different aspect suggests that the SemCSE-Multi can effectively disentangle and isolate aspect-specific information.
@@ 4.3 Results
Results are displayed in Table 2. The upper part reports the correlations between the aspect-specific embeddings created by SemCSE-Multi and the different ground-truth similarity assessments. For completeness, we include the human-annotated hypothesis labels as predictors (row 1), allowing us to assess their correlation with the LLM-produced similarity scores. We see that the human-annotated hypothesis labels show high correlation with the LLM-annotated hypothesis similarity scores, thus demonstrating that the LLM-based annotation is a reasonable approach.
@@
    As discussed, the human-annotated hypothesis labels strongly correlate with the LLM-based similarities for the hypothesis aspect.
@@
    Embeddings for the hypothesis aspect correlate well with the human-annotated hypothesis labels, thus demonstrating that this factor is effectively represented in the embeddings.
@@
The substantially lower correlations observed in the remaining off-diagonal cells provide strong and statistically significant (Appendix A.6) evidence that the aspect-specific embeddings are indeed isolating the intended factors. In consequence, this disentanglement enables controlled similarity assessments adhering to specific user needs.
One notable exception is the recommendation aspect, where embeddings only show weak alignment with the LLM-based similarities. This is likely due to the scarcity of explicit recommendations in scientific abstracts, which reduces the number of training samples, and also reduces data quality due to the increased the likelihood of the LLM hallucinating summaries if recommendations are absent or solely hinted at within the abstract.
To contextualize the strength of the observed correlations, we additionally evaluate several widely used scientific embedding models: SPECTER Cohan et al. (2020), SPECTER2 Proximity Singh et al. (2023), SciNCL Ostendorff et al. (2022b), SemCSE Brinner and Zarriess (2025), and the invasion biology-specific InvDef-DeBERTa Brinner et al. (2025). We do not include other multifaceted models, as (to our knowledge) none of them allow for freely selecting the target aspects without access to corresponding supervised datasets.
Across all aspects except recommendation, none of these models achieve correlations comparable to our aspect-specific embeddings, thus demonstrating the benefits of having an aspect-specific focus. More importantly, these baseline models cannot disentangle specific aspects, making them unsuitable for tasks requiring fine-grained similarity assessments. This is underscored by Figure 1, visualizing the clearly differing and semantically structured aspect-specific embedding spaces by SemCSE-Multi for the Hypothesis and Species aspects, as well as the substantially less distinctly structured space induced by the SciNCL model. Together, these results provide clear evidence that a multifaceted embedding model offers distinct advantages both in terms of accuracy and in the ability to isolate aspect-specific information.
We further find that using a two-stage procedure is crucial for disentangling the individual aspects. We evaluated an additional baseline (Appendix A.7) that directly trains a unified model for all aspects, but find that correlations between different aspects are substantially higher, likely caused by the co-adaption of the model to all aspects at once, which leaks information between them. Training aspect-specific models in isolation and distilling a unified model to predict these properly isolated embeddings is therefore a crucial advantage.
@@ 5 Embedding Decoding
In this section, we demonstrate that it is possible to use a decoder LLM to map aspect-specific embeddings back into the space of natural language descriptions, thus substantially enhancing the interpretability of the embedding space.
@@ 5.1 Model Training
To guide the decoding process, we additionally prepend five trainable, input-independent embeddings that function as general task descriptors, and optionally append one more token to act as delimiter to the decoded text, in case a normal autoregressive language model is used instead of a chat model. During training, only these embeddings and the mapping network are optimized, while the parameters of the underlying LLM remain frozen. Using the dataset of summarizing sentences paired with their corresponding aspect-specific embeddings, the model is trained with a standard language modeling cross-entropy loss to reconstruct the original summary from its embedding representation.
We evaluate two LLM backbones of different sizes and training paradigms, with one being a purely autoregressive language model in the form of Llama 3 8B Grattafiori et al. (2024), and the other being a larger (24B parameters) chat model in the form of Mistral Small 3.1 Mistral AI (2025).
@@ 5.2 Evaluation Setup
BERTScore
Hypotheses general 77.5 76.1 71.9 - 79.2 77.6 72.9
Hypotheses specific 76.5 75.5 71.6 - 78.0 76.5 71.2
Ecosystem 77.2 76.1 73.0 - 78.6 77.4 73.7
Res. Question 74.5 72.3 69.0 - 75.9 74.1 70.4
Species 75.5 74.5 68.8 - 76.8 75.5 69.7
Methodology 78.7 77.8 72.5 - 79.9 78.9 72.9
Recommendation 70.1 69.4 67.8 - 70.6 69.0 68.5
Table 3: Perplexity and BERTScore results for Llama and Mistral decoders when decoding embeddings under normal, reconstructed and randomized settings. Lower perplexity indicates better performance, while higher BERTScores indicate better alignment of generated descriptions with ground-truth descriptions.
We evaluate the decoding pipeline on the test set of 954 abstracts for which we predict aspect embeddings with SemCSE-Multi. For a given aspect embedding, we first measure the perplexity of the ground-truth summaries under the predicted distribution produced by the LLM, so that low perplexity indicates that the decoder can recover the semantic content of the original summary with high probability. As a second evaluation metric evaluating actual generated sentences, we use BERTScore Zhang et al. (2020) on a pretrained SciBERT model Beltagy et al. (2019), which evaluates the correspondence between generated and ground-truth tokens.
To verify that the model does not merely generate generic aspect-related sentences that achieve reasonable performance while ignoring the actual embedding input, we additionally evaluate both metrics when matching the embedding of one sample with the summaries of another (Shuffled), and, for the pure autoregressive Llama model, without any embedding information (Unconditioned). If the trained decoding process indeed reconstructs the semantic meaning of the embeddings, using matching embedding and ground-truth should result in substantially lower perplexity and higher BERTScores compared to the control experiments.